if jan is 12 years younger than Peter. After 2 years Peter will be three times older than Jan then what is the age of both now _ make a equation .

Answer by addingup(3677)

P = J+12
P+2 = 3(J+2)
There are your two equations. Distribute the 3 over the terms inside the parenthesis on the right, substitute for P on the left, and apply algebra to solve. Alwas do to one side what you do to the other.